You’ve observed the following returns on Crash-n-Burn Computer’s stock over the past five years: 10 percent, –10 percent, 17 percent, 22 percent, and 10 percent.

a.

What was the arithmetic average return on Crash-n-Burn’s stock over this five-year period? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place, e.g., 32.1.)

  Average return %
b-1

What was the variance of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations and round your answer to 5 decimal places, e.g., 32.16161.)

Variance  

b-2

What was the standard deviation of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Standard deviation %
Answer #1

Average return=Total return/Total time period

=(10-10+17+22+10)/5=9.8%

Return (Return-Average return)^2
10 (10-9.8)^2=0.04
-10 (-10-9.8)^2=392.04
17 (17-9.8)^2=51.84
22 (22-9.8)^2=148.84
10 (10-9.8)^2=0.04
Total=592.8%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[592.8/(5-1)]^(1/2)

=12.17%(Approx).

Variance=Standard deviation^2

=0.01482